OpenStudy (anonymous):

55 mi / h = ____ ft / s

OpenStudy (anonymous):

290000

OpenStudy (anonymous):

approx.

OpenStudy (anonymous):

I'm too lazy to do this, so I'll use google! Google tells me it's 80.6 ft/s :)

OpenStudy (anonymous):

The amount of feet cannot be fewer than the amount of miles.

OpenStudy (anonymous):

It's asking feet per second, you do realize that right? If that person was traveling at 290000 feet per second, it would be 19772.7272727 miles per hour.

OpenStudy (anonymous):

With what you said 290,000 feet/ second...that's 197,727 miles/hr...

OpenStudy (anonymous):

That's some insane speed a person or w e is moving :O

OpenStudy (anonymous):

\[\frac{ 55 mph }{ 1 } \times \frac{ 1 hr }{ 60\min } \times \frac{ 1\min }{ 60 \sec } \times \frac{ 5280 ft }{ mile } = \frac{ 55 \times 5280 }{ 60 \times 60 }\ = \frac{ 290400 }{ 3600 } = 80.666666 ft/s\]]